A fairgrounds ride spins its occupants inside a flying-saucer-shaped container. If the horizontal circular path the riders follow has a 5.00 m radius, at how many revolutions per minute will the riders be subjected to a centripetal acceleration 2.65 times that of gravity? x* rev/min
Answer #1

Solution :-

W = mg

Fc = mV²/R

Fc = 2.65(mg)

mV²/R = 2.65mg

V²/R = 2.65 g

V² = 2.65 Rg

V = √2.65(5)(9.8)

V = 11.4 ms

wR = V

w = V/R

w = 11.4/5= 2.28 rad/s

w = 2.28(60)/ 2π

w = 21.78 rev/min